undeadknight26 (undeadknight26):

The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 480 feet. Show all steps in your work.

undeadknight26 (undeadknight26):

I have to solve d = .5gt2

undeadknight26 (undeadknight26):

@ganeshie8 @phi @thomaster @radar @dan815 @Loser66 please help!

undeadknight26 (undeadknight26):

@michelle_DrNiah @marylou004 @hihi67 @Xmoses1 @Niall

undeadknight26 (undeadknight26):

@Luigi0210 please help....

OpenStudy (phi):

*** Solve this equation for g, *** \[ d = \frac{1}{2}gt^2 \] what do you get (as a first step) when you multiply both sides by 2 ?

undeadknight26 (undeadknight26):

2d = 1gt^2?

OpenStudy (phi):

yes. People don't bother to show 1 * anything because multiplying by 1 doesn't "do" anything: \[ 2d = g t^2 \] now divide both sides by t^2

undeadknight26 (undeadknight26):

2d = gt^2 /t^2 /t^2 2d = g?

OpenStudy (phi):

the right side is ok because t^2/ t^2 is 1 (anything divided by itself is one) what happened to t^2 on the left side? It should be there

OpenStudy (phi):

in other words, the left side is 2d divided by t^2

OpenStudy (phi):

can you write down your equation for g ?

undeadknight26 (undeadknight26):

sorry i was afk...i do not undeadstand...is it d = g?

OpenStudy (phi):

no. Dividing in algebra is easy. you put /t^2 next to 2d \[ 2d = g t^2 \] divide both sides by t^2 \[ \frac{2d }{t^2} = \frac{gt^2}{t^2} \] now simplify (if you can)

undeadknight26 (undeadknight26):

You cannot simplify this equation right? or if you could it would be d/t = g/t

OpenStudy (phi):

what about the right side ? you know that \[ \frac{gt^2}{t^2} = g \frac{t^2}{t^2} \]

undeadknight26 (undeadknight26):

so d = g t^2/t^2?

OpenStudy (phi):

\[ \frac{2d }{t^2} = \frac{gt^2}{t^2} \] do you see that t^2 does not divide into 2 or d ? (on the left side) on the right side, t^2 does not divide into g, but it does divide into t^2

undeadknight26 (undeadknight26):

Yes. So does that mean the 2d/t^2 = gt^2/t^2 is my final answer?

OpenStudy (phi):

I think you are missing the point. 1) you are solving for g. that means g *by itself* on one side of the equation 2) t^2/t^2 simplifies to 1